OpenStudy (anonymous):

A bag contains 50 marbles, 10 of which are blue, 8 are red, 20 are green, and 12 are purple. Marcie takes a marble out of the bag, records the color, and returns it to the bag. How many red marbles should she expect after 500 trials?

OpenStudy (muscrat123):

hang on :)

OpenStudy (anonymous):

ok :)

OpenStudy (muscrat123):

What is the faction of red marbles in the bag?

OpenStudy (muscrat123):

# of red / total number of marbles what do you get ?

OpenStudy (muscrat123):

10 of which are blue, 8 are red, 20 are green, and 12 are purple Checking: 10 + 8 + 20 + 12 = 50 -- Good Theoretical Ratios: Blue 10/50 = 20%, Red 8/50 = 16%, Green 20/50 = 40%, Purple 12/50 = 24% Checking: 20 + 16 + 40 + 24 = 100 -- Good!

OpenStudy (anonymous):

wait so my answer is 100? @muscrat123

OpenStudy (muscrat123):

i think so, r there choices?